Circular No. 7229

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999dg IN UGC 9758 M. Modjaz and W. D. Li, University of California at Berkeley, on

behalf of the Lick Observatory Supernova Search (cf. IAUC 6627, 7126), report the discovery with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) of an apparent supernova in UGC 9758 (= MCG +02-39-007). SN 1999dg was discovered on unfiltered images taken on July 23.2 UT (mag ∼ 16.7) and confirmed on images taken on July 24.2 (mag ∼ 16.5). The new object is located at α = 15^{h}11^{m}28^{s}.41, δ = +13^{o}29′05′′.7 (equinox 2000.0), which is 4′′.8 east and 5′′.2 north of the nucleus of UGC 9758. A KAIT image of the same field on July 15.2 (limiting mag ∼ 19.0) showed nothing at the position of SN 1999df.

SUPERNOVA 1999dh IN IC 211 Modjaz and Li also report the discovery (as above) of an apparent

supernova in IC 211. SN 1999dh was discovered on unfiltered images taken on July 23.5 UT (mag ∼ 15.4) and confirmed on images taken on July 24.5 (mag ∼ 15.6). The new object is located at α = 2^{h}11^{m}06^{s}.07, δ = +3^{o}51′06′′.9 (equinox 2000.0), which is 29′′.0 west and 1′′.7 south of the nucleus of IC 211. A KAIT image of the same field on Feb. 4.2 (limiting mag ∼ 19.0) showed nothing at the position of SN 1999dh.

COMET C/1999 N2 (LYNN) Total visual magnitude and coma-diameter estimates: July 15.87 UT,

7.7, 6′ (J. G. de S. Aguiar, Joaquim Egidio, Brazil, 0.23-m reflector); 16.39, 7.0, 5′ (M. Mattiazzo, Wallaroo, South Australia, 7×50 binoculars); 17.26, 6.8, 5′ (N. Biver, Oahu, HI, 7×50 binoculars); 18.37, 7.1, 7′ (C. E. Drescher, Beenleigh, Qld., Australia, 20×80 binoculars); 20.34, 7.1, 4′ (D. A. J. Seargent, The Entrance, N.S.W., 10×50 binoculars); 22.46, 7.3, 5′.5 (A. Pearce, Nedlands, W. Australia, 20×80 binoculars); 23.46, 7.2, 7′ (Pearce).

COMET C/1999 J3 (LINEAR) Visual m_{1} estimates: May 15.98 UT, 15.2 (R. J. Bouma, Lattrop, The

Netherlands, 0.40-m reflector); July 10.97, 14.0 (Bouma); 12.17, 12.7 (A. Hale, Cloudcroft, NM, 0.41-m reflector); 16.28, 12.2 (N. Biver, Oahu, HI, 0.26-m reflector); 17.95, 11.4 (M. Meyer, Frauenstein, Germany, 0.25-m reflector); 21.97, 11.9 (M. Reszelski, Szamotuly, Poland, 0.25-m reflector).

1999 July 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green